# Topological Solitons in Su-Schrieffer-Heeger Chain with periodic hopping modulation, domain walls and disorder

### v1

@@ Abstract
A chiral symmetric Su-Schrieffer-Heeger (SSH) chain features topological end states in one of its dimerized configurations. Those mid-gap zero energy states show interesting modifications upon a periodic tuning of the hopping modulations. Besides, more and more in-gap end modes appear at nonzero energies for further partitioning of the Brillouin zone due to increased hopping periodicity. The new topological phases are identified with a detailed analysis of the topological invariants namely, winding number and Zak phases. Spectra and topology of these systems with periodically modulated hopping are studied also in presence of a single static domain wall, separating two topologically inequivalent dimerized structures. Lastly, we also study the effect of disorder, particularly the chirality breaking onsite ones, on the edge and domain wall states. The end solitons and domain wall solitons show noteworthy evolutions with the variation of hopping periodicity. Our findings can add important feedback in utilizing topological phases in various fields including quantum computations and the results can be easily verified in a cold atom set up within optical lattices.
@@ I Introduction
One can mention of graphene in the journey of topological condensed matter which garnered huge attention after its discovery[17] using scotch tape method in 2004 due to its Dirac like excitations at low energies, high mobilities and topological stability[18]. From there on, physicists have moved on a lot via Haldane model[19], Kane-Mele model[20], topological insulators[21] and then Weyl, Dirac semimetals[22] and so forth. Graphene has a staggered single bond-double bond structures (considering single resonance structure[23]) in a hexagonal lattice. But before it became popular, scientists were intrigued by the 1D conducting polymers involving same staggered bonding structures. Thats when the $(CH)_{n}$ Polyacetylene chain[24] became important to the science community for its topological behavior, solitonic excitations and domain wall structures[25, 26, 27].
@@
A long chain of Polyacetylene has a pair of degenerate ground states for two different set of dimerization[25]. For a finite chain these two staggered arrangements are topologically distinct for the end site sees either a strong or a weak bond. In the topological regime, end modes appear as zero energy states (ZES) that peak at the boundaries and die out exponentially away from there. One can call those end solitons. In a half filled system, their spectral weights are equally shared between unoccupied conduction and occupied valence bands. With electron added to or removed from the half filled system, the end modes occupy with fractionalized charges $\mp e/2$. Strictly speaking, they are bonding and antibonding small gap states and of mixed chirality. But linear combinations of them can be considered to obtain a pair of chiral ZES that are located in single ends of the chain[28].
@@
A Su-Schrieffer-Heeger (SSH) model that can correspond to a polyacetylene chain, is a 1D tight binding model with staggered hopping modulation[25] and it can demonstrate charge fractionalizations[29, 30], the existence of zero energy end states and topological solitonic excitations[25, 27]. Neural solitons with moving domain walls having $S=\pm\frac{1}{2}$ are obtained as the excitations though charged solitons can also found for a doped $(CH)_{n}$ system[26]. The SSH chain has chiral symmetry and the chain features unit cells comprising of two adjacent sites. The staggered nature of the hopping brings in this two sublattice structure in the chain and accordingly the chiral symmetry, here, is also familiar as sublattice symmetry. It makes the Bloch Hamiltonian off-diagonal which gives the simplest route to derive the topological winding number, as we also have elaborated in this paper.
@@
In this respect it will not be ludicrous to add a discussion on exotic Majorana fermions, that are their own antiparticles which appear as quasiparicle excitations (called Majorana bound states) in some condensed matter systems with defects[1, 2, 3]. Our system of concern - the SSH model can be broken down to two independent Majorana hopping chains. Similarly in a spinless Kitaev chain with $p$-wave pairing[31], a pair of independent Majorana hopping chains appear and single Majorana zero modes (MZM) survive there at the ends. The unconvensional superconducting pairing can nullify the effect of quantum zero point motion to pin the Majorana states at the zero energy. Due to their non abelian exchange statistics, they lead to decoherence-free quantum information processing[14]. However for the SSH model, the pair of Majorana modes localized at each end turn them into electronic modes and we do not get any Majorana physics in this system32. But with interaction added, Majorana modes can be incorporated in such systems.
@@
The chiral symmetric SSH chain, with Hamiltonian as shown in Eq.(1), features a topological phase for weak bonds at the boundaries and midgap zero energy states (ZES) with fractional charge are obtained. Fig. 1 demonstrates that while in a Kitaev chain end states are found to be Majorana modes, in SSH model pair of Majorana modes at each end produces electronic ZES at two boundaries. Interesting becomes the case where ZES gets redistributed/modified due to a periodic variation of the hopping modulation that goes beyond simple staggered hopping of the SSH chain. Similar study has been initiated by one of the authors in Ref.12. In this paper, we extend the calculations incorporating detailed analysis of the winding numbers to identify the topological/non-topological end modes of this periodically hopping modulated system. In addition here we also consider the effect of domain wall and disorders. The consequence having domain walls, in the SSH chain[34] and other related models[26, 35, 36, 37] have been probed to some extent in the literature and hence our attempt to study them in presence of periodic modulation of hopping amplitude will boost further understanding of the topic.
@@ II Formulation
The Su-Schrieffer-Heeger (SSH) model [25], proposed in the context of polyacetylene is given by a one-dimensional tight-binding Hamiltonian with staggered nearest neighbor hopping (thus it shows a chiral (sub-lattice) symmetry:
@@
here $c_{i}^{\dagger}$, $c_{i}$ denotes electron creation and annihilation operator respectively and the periodic modulation in hopping strength (t) is obtained by $\delta_{i}=\Delta cos[(i-1)\theta]$ in which $i=1,2,3,......,N$. In general, for $\theta=\frac{2\pi}{n}$, we get $\delta_{i+1}=\Delta cos(\frac{2\pi i}{n})$ and the chain froms a $n$ sublattice structure. Moreover, the system is represented by a $n\times n$ Hamiltonian matrix with $n$ number of eigenmodes. The transformation $c_{i}\rightarrow(-1)^{i}c_{i}$ gives $H_{SSH}\rightarrow-H_{SSH}$, which express the chiral symmetry. For a model considering periodic boundary conditions, chiral symmetry also needs the total number of sites to be even.
@@ II.1 For $\theta=\frac{\pi}{2}$
where $i$ denotes the unit cell. For periodic boundary condition (PBC), we can consider the Fourier transform for the above Hamiltonian as
$c_{i,\alpha}=\frac{2}{\sqrt{L}}\sum_{k}e^{jki}\beta$ (5)
where $\alpha=\{A,B,C,D\}$ and $\beta=\{a_{k},b_{k},c_{k},d_{k}\}$ refer to the sublattice index and the corresponding Fourier modes respectively. Thus the Bloch Hamiltonian takes the form of
$H$ $=$ $\sum_{k}\Bigg{[}(t+\Delta)(a_{k}^{\dagger}b_{k}+b_{k}^{\dagger}a_{k})+t(b_{k}^{\dagger}c_{k}+c_{k}^{\dagger}b_{k})+$ (6)
$(t-\Delta)(c_{k}^{\dagger}d_{k}+d_{k}^{\dagger}c_{k})+t(d_{k}^{\dagger}a_{k}e^{ik}+a_{k}^{\dagger}d_{k}e^{-ik})\Bigg{]}$
$=$ $\sum_{k}\psi_{k}^{\dagger}H_{k}\psi_{k}$
where
$H_{k}=\begin{pmatrix}0&(t+\Delta)&0&te^{-ik}\\ (t+\Delta)&0&t&0\\ 0&t&0&(t-\Delta)\\ te^{ik}&0&(t-\Delta)&0\end{pmatrix}$ (7)
and
$\psi_{k}=(a_{k},b_{k},c_{k},d_{k})^{T}$
@@
$\epsilon(k)=\pm\sqrt{2t^{2}+\Delta^{2}\pm 2t\sqrt{2\Delta^{2}+(t^{2}-\Delta^{2})cos^{2}k/2}}.$ (8)
It essentially gives four bands in the energy spectrum. Notice that the energy gap closes for $\Delta=0$ at $k=0$ as well as for $(\Delta/t)^{2}=2$ at $k=\pm\pi$. Of these, the 1st case is a result of using periodic boundary implied in our Fourier construction but the system gaps out there in a finite chain. However the 2nd case indicates a gapless point even for the finite chain and correspond to the gap closing topological phase transition point. The bands cross linearly there (see Fig.3) making the low energy modes behave as Dirac fermions[38].
The numerical energy spectra for SSH(like) chain under OBC for $\theta=\pi/2$ is shown in Fig.2. The figure indicates that energy spectra is symmetric (up-down) i.e. for every eigenstate with energy $E$, its chiral symmetric partner having energy $-E$ necessarily exists. Here, we get a 4 sublattice configuration yielding 2 in-gap non-zero energy states and 2 ZES. Zero energy modes with $L=256$ will no longer subsist for $-0.3\lesssim\Delta/t\lesssim 0.3$ and for $\Delta/t\gtrsim 1.4$ or $\Delta/t\gtrsim-1.4$ ( see e.g. Figs.2-2).
For an odd number of lattice sites, one gets odd number of zero energy states. We examined that in a chain with odd number of sites, there are $L~{}modulo~{}4$ number of extra ZES that are localized exactly at $Mod(L,4)$ end sites. For example for $L=255$, with spectra as shownin in Fig.2, there are 5 (for $|\Delta/t|\gtrsim 0.25$) or 3 (for $|\Delta/t|\lesssim 0.25$) zero energy modes ($i.e.$, 3 extra ZES) with the extra modes being localized at site no. 253, 254 and 255 (it could have been at site number 1, 2, 3 on the other side of the chain had the Hamiltonian matrix taken the three extra bonds on that side).
Figure 2: Spectra of a SSH(like) chain considering OBC as a function of $\Delta/t$ for $\theta=\frac{\pi}{2}$. (b) and (c) shows the corresponding low energy states with smaller and larger $\Delta/t$ respectively while (d) represents low energy spectra for an odd number of lattice sites.
Class
T
C
S
d=1
d=2
d=3
A
0
0
0
Z
AI
+1
0
0
AII
-1
0
0
$\textbf{Z}_{2}$
$\textbf{Z}_{2}$
D
0
+1
0
$\textbf{Z}_{2}$
Z
C
0
-1
0
Z
AIII
0
0
1
Z
Z
BDI
+1
+1
1
Z
CI
+1
-1
1
Z
DIII
-1
+1
1
$\textbf{Z}_{2}$
$\textbf{Z}_{2}$
Z
CII
-1
-1
1
Z
$\textbf{Z}_{2}$
@@ II.1.1 Symmetry and Winding Number
Now we discuss a bit more on the symmetry of the SSH(like) chain for $\theta=\frac{\pi}{2}$. By employing the matrix $\Gamma_{4}=I_{2}\otimes\sigma_{z}$ with $I_{2}$ is the $2\times 2$ identity matrix, one can check that $\{\Gamma_{4},H_{k}\}=0$. So, $\Gamma_{4}$ is the chiral operator that anti-commutes with $H_{K}$. Hence, this Hamiltonian has a chiral symmetry (of course, for a chain with even number of sites). Apart from this, this Hamiltonian also respects the time-reversal symmetry ($T$) as well as particle-hole symmetry ($C$) and falls in the BDI class universality in a 10-fold way classification (see Table 1). Due to the chiral symmetry of Hamiltonian, the Bloch Hamiltonian (Eq.(7)) is expected to be off-diagonal. We should add here that the presence of chiral symmetry ($S$) does not ncessarily ensure $T$ and $C$ symmetry as there is a symmetry class called AIII that respects chiral symmetry alone. In fact, if we consider the hopping amplitudes to be complex, instead of real, then time-reversal symmetry no longer persists and $H_{SSH}$ moves on to the AIII universality class[41]. The AIII class, in one dimension ($1D$), is always topologically nontrivial[39], unlike the usual SSH chain (with $\theta=\pi$), and its topology is described by a topological winding number.
There are two important corollaries of this chiral symmetry for the eigenstates and energies of the Hamiltonian. Firstly, the energy spectrum of the SSH(like) chain is always symmetric (up-down symmetry) in nature. For each eigenstate [$\psi=(\psi_{A},\psi_{B})^{T}$] having energy $E$, there is always another eigenstate [its chiral partner, $\Gamma\psi=\psi^{\prime}=(\psi_{A},-\psi_{B})^{T}$] with energy $-E$. Secondly, states with zero energy only occupy a single sub-lattice. When the total number of particles is odd, chiral symmetry inferred that there will be an odd number of end modes having zero energy. Moreover, the energy of the in-gap end states can also be non-zero, and this symmetry implies that the such states having non-zero energy should be even in number and symmetrically positioned about $E=0$.
@@
$V(k)=\begin{pmatrix}(t+\Delta)&te^{-ik}\\ t&(t-\Delta)\end{pmatrix}$ (10)
@@
Now, the calculation of the winding number can be given by [39, 43, 48, 44, 47]
@@
with integration on the reduced BZ, $k\in$ [$-\pi$$\pi$] and the estimate comes out to be
$\mathcal{W}=\Bigg{\{}\begin{matrix}1,&|t^{2}-\Delta^{2}|<|t^{2}|;\\ 0,&|t^{2}-\Delta^{2}|>|t^{2}|;\\ \text{undefined},&|t^{2}-\Delta^{2}|=|t^{2}|\end{matrix}$ (13)
The complex variable Det[V(k)]=$(t^{2}-\Delta^{2})-t^{2}e^{-ik}$ is called the winding vector (it has a magnitude and phase angle in the complex plane) as its number of winding about the origin as $k$ varies in [$-\pi$, $\pi$] [49] determines the winding number.
Eq.13 implies that, one can get $\mathcal{W}=1$ for small $\Delta/t\neq 0$ indicating the non-trivial topological phases (NTP) there. But $\Delta/t=0$ indicates an undefined ${\mathcal{W}}$ and hence no gap closing Lifshitz quantum phase transition (QPT) point. On the other hand, ${\mathcal{W}}$ remains unity till $0<|\Delta/t|<\sqrt{2}$ and vanishes for $|\Delta/t|>\sqrt{2}$. Thus QPT occurs at $\Delta=\pm\sqrt{2}t$ in this case. We should add here that interchanging the Majorana operators $\gamma_{i,A}$, $\gamma_{i,B}$ in Hamiltonian Eq.(3) has no significant effect (other than an overall sign change) since our considered model is not including the next-nearest-neighbor hopping amplitudes and chemical potentials. Following the above procedure, we attain
$V^{\prime}(k)=\begin{pmatrix}(t+\Delta)&t\\ te^{ik}&(t-\Delta)\end{pmatrix}$ (14)
Now, the winding number is
$\mathcal{W}=\Bigg{\{}\begin{matrix}-1,&|t^{2}-\Delta^{2}|<|t^{2}|;\\ 0,&|t^{2}-\Delta^{2}|>|t^{2}|;\\ \text{undefined},&|t^{2}-\Delta^{2}|=|t^{2}|\end{matrix}$ (15)
It is worth mentioning that the winding number changes sign ($\mathcal{W}\rightarrow-\mathcal{W}$) by relabelling sublattice $A\Leftrightarrow B$(or, interchanging the Majorana operators). This can also be visualized as a ‘choice’ of chiral symmetry operators. In the former case it is $S=\Gamma_{4}$ acting in the sublattice basis while in the latter it is $S=-\Gamma_{4}$. Moreover, the sign of the winding number is related to the winding direction of winding vector Det[V(k)] which in turn leads to the type of (unpaired) Majorana fermion residing at the boundary. This change of sign makes no difference when one considers a single chain, however for multiple chains coupled to each other, such liberty no more persists as the relative sign of the individual chains become a relavant quantity of the system topology and Hamiltonian no more remains block off-diagonal. Detailed illustration of the same can be found in Ref.[42].
@@ II.1.2 Energy-Band analysis and Beyond
As already mentioned, one can write
${\bf Det[V(k)]}=r_{1}(k)+ir_{2}(k)=R(k)e^{i\phi(k)}$ (16)
where $\phi(k)$ is the phase of Det[V(k)] in the complex plane and $r_{1}$, $r_{2}$ are real numbers. This leads to[42, 46]
$\mathcal{W}=-\frac{i}{2\pi}\int_{BZ}\partial_{k}i\phi(k)dk\ .$ (17)
For a two band model (see e.g. Refs.[45, 50]) one does not need to consider the determinant as $V(k)$ is just a number. However for systems with 4 bands or more, this formula aims to give the topological index without further diagonalizing the Hamiltonian Eq.(7). Consequently, Eq.(17) takes the form of
$\mathcal{W}=\frac{1}{2\pi}[\phi(\pi)-\phi(-\pi)]=\frac{1}{2\pi}\Delta\phi$ (18)

### v2

# Topological Solitons in Su-Schrieffer-Heeger Chain with periodic hopping modulation, domain wall and disorder

@@ Abstract
A chiral symmetric Su-Schrieffer-Heeger (SSH) chain features topological end states in one of its dimerized configurations. Those mid-gap zero energy states show interesting modifications upon a periodic tuning of the hopping modulations. Besides, more and more in-gap end modes appear at nonzero energies for further partitioning of the Brillouin zone (BZ) due to increased hopping periodicity. The new topological phases are identified with a detailed analysis of the topological invariants namely, winding number and Zak phases. The spectra and topology of these systems with periodically modulated hopping are studied also in the presence of a single static domain wall, separating two topologically inequivalent dimerized structures. The domain wall causes additional in-gap modes in the spectrum as well as zero energy domain wall solitonic states for specific hopping periodicities. We also study the effect of disorder, particularly the chirality breaking onsite ones, on the edge and domain wall states. Other than the SSH type we also consider random, Rice-Mele or AI type disorder to do a comparative analysis of the evolution of chirality and zero energy states as the strength of disorder and hopping periodicity is varied. Our findings can add important feedback in utilizing topological phases in various fields including quantum computations while the results can be easily verified in a cold atom set up within optical lattices.
@@ I Introduction
Today we often talk of graphene whose discovery[17] using scotch tape method in 2004 turned out a milestone in the journey of topological condensed matter systems, for its Dirac like excitations at low energies, high mobilities and topological stability[18]. From there on, physicists have moved on a lot on topological system synthesis/analysis via introducing Haldane model[19], Kane-Mele model[20], topological insulators[21] and then Weyl, Dirac semimetals[22] and so forth. Graphene has a staggered single bond-double bond structure (considering single resonance structure[23]) in a hexagonal lattice. But before it became popular, scientists were intrigued by the 1D conducting polymers involving the same staggered bonding structures. That’s when the $(CH)_{n}$ Polyacetylene chain[24] became important to the science community for its topological behavior, solitonic excitations, and domain wall structures[25, 26, 27].
@@
A long chain of Polyacetylene has a pair of degenerate ground states for two different sets of dimerization[25]. For a finite chain these two staggered arrangements are topologically distinct for the end site sees either a strong or a weak bond. In the topological regime, end modes appear as zero energy states (ZES) that peak at the boundaries and die out exponentially away from there. One can call those end solitons. In a half-filled system, their spectral weights are equally shared between unoccupied conduction and occupied valence bands. With electrons added to or removed from the half-filled system, the end modes occupy with fractionalized charges $\mp e/2$. Strictly speaking, they are bonding and antibonding small gap states and of mixed chirality. But linear combinations of them can be considered to obtain a pair of chiral ZES that are located in single ends of the chain[28].
@@
A Su-Schrieffer-Heeger (SSH) model that can correspond to a polyacetylene chain, is a 1D tight-binding model with staggered hopping modulation[25] and it can demonstrate charge fractionalizations[29, 30], the existence of zero energy end states and topological solitonic excitations[25, 27]. Neutral solitons with moving domain walls having $S=\pm\frac{1}{2}$ are obtained as the excitations though charged solitons can also found for a doped $(CH)_{n}$ system[26]. The SSH chain has chiral symmetry and the chain features unit cells comprising of two adjacent sites. The staggered nature of the hopping brings in this two sublattice structure in the chain and accordingly the chiral symmetry, here, is also familiar as sublattice symmetry. It makes the Bloch Hamiltonian off-diagonal which gives the simplest route to derive the topological winding number, as we also have elaborated in this paper.
@@
In this respect it will not be ludicrous to add a discussion on exotic Majorana fermions, that are their own antiparticles which appear as quasiparicle excitations, called Majorana bound states (MBS) in some condensed matter systems with defects[1, 2, 3]. Our system of concern - the SSH model can be broken down to two independent Majorana hopping chains. Similarly in a spinless Kitaev chain with $p$-wave pairing[31], a pair of independent Majorana hopping chains appear and single Majorana zero modes (MZM) survive there at the ends. The unconventional superconducting pairing can nullify the effect of quantum zero point motion to pin the Majorana states at the zero energy. Due to their non abelian exchange statistics, they lead to decoherence-free quantum information processing[14]. However for the SSH model, the pair of Majorana modes localized at each end turn them into electronic modes and we do not get any Majorana physics in this system[32]. But with interaction added, Majorana modes can be incorporated in such systems.
Refer to caption(a)
Refer to caption(b)
@@
The chiral symmetric SSH chain, with Hamiltonian as shown in Eq.(1), features a topological phase for weak bonds at the boundaries, and midgap zero energy states (ZES) with fractional charge are obtained. Fig.1 demonstrates that while in a Kitaev chain end states are found to be Majorana modes, in the SSH model pair of Majorana modes at each end produces electronic ZES at two boundaries. Interesting becomes the case where ZES gets redistributed/modified due to a periodic variation of the hopping modulation that goes beyond simple staggered hopping of the SSH chain. A similar study has been initiated by one of the authors in Ref.12. In this paper, we extend the calculations incorporating detailed analysis of the winding numbers to identify the topological/non-topological end modes of this periodically hopping modulated system. In addition here we also consider the effect of domain wall (DW) and disorders. A static DW at the center of a SSH chain break it into two topologically inequivalent regime on its two sides and results in solitonic states in the system. The consequence of having domain walls, in the SSH chain[34] and other related models[26, 35, 36, 37] have been probed to some extent in the literature. Here we find that the presence of periodic modulation of hopping amplitude causes additional in-gap modes to appear. Besides one of the end modes vanishes with new zero energy solitonic states appearing at the DW position for specific periodicities of hopping modulations. Furthermore, the addition of onsite disorder in our model leads to immediate disappearance of the chiral symmetry. We also consider random[38], Rice-Mele (RM) and AI-type of disorder[39] for comparison which reveals that the disappearance of ZES with disorder strength is common in all of them though a DW state approaches zero energy for strong disorder of AI-type.
The paper is organized as follows. In Sec.II we provide the formulation of our SSH(like) model with periodic hopping modulation including also a detailed analysis of symmetry, topological invariant calculations as well as the features of spectra and end states. Next in Sec.III we additionally introduce a single static domain wall at the center of a finite chain and study its response on the ZES and energy spectra. Sec.IV describes the effect of onsite and hopping disorder in the periodically modulated chain with domain wall which affects the chirality of the system. Finally, we summarize our results in Sec.V and discuss its novelty and possible future directions of work.
@@ II Formulation
The Su-Schrieffer-Heeger (SSH) model [25], proposed in the context of polyacetylene is given by a one-dimensional tight-binding Hamiltonian with staggered nearest neighbor hopping [thus it shows a chiral (sub-lattice) symmetry]:
@@
here $c_{i}^{\dagger}$, $c_{i}$ denotes electron creation and annihilation operator respectively and the periodic modulation in hopping strength (t) is obtained by $\delta_{i}=\Delta\cos[(i-1)\theta]$ in which $i=1,2,3,......,N$. In general, for $\theta=2\pi/n$, we get $\delta_{i+1}=\Delta\cos(\frac{2\pi i}{n})$ and the chain from a $n$ sublattice structure. Moreover, the system is represented by a $n\times n$ Hamiltonian matrix with $n$ number of eigenmodes. The transformation $c_{i}\rightarrow(-1)^{i}c_{i}^{\dagger}$ gives $H_{SSH}\rightarrow H_{SSH}$, implying the sublattice or chiral symmetry[40]. For a model considering periodic boundary conditions (PBC), chiral symmetry also needs the total number of sites to be even.
@@ II.1 For $\theta=\frac{\pi}{2}$
where $i$ denotes the unit cell. For PBC, we can consider the Fourier transform for the above Hamiltonian as
$c_{i,\alpha}=\frac{2}{\sqrt{L}}\sum_{{k\epsilon BZ}}e^{jki}\beta$ (5)
where $\alpha=\{A,B,C,D\}$ and $\beta=\{a_{k},b_{k},c_{k},d_{k}\}$ refer to the sublattice index and the corresponding Fourier modes respectively. Thus the Hamiltonian [Eq.(4)] in terms of the spinor field $\psi_{k}=(a_{k},b_{k},c_{k},d_{k})^{T}$ takes the form of
$H$ $=$ $\sum_{{k\epsilon BZ}}\psi_{k}^{\dagger}H_{k}\psi_{k}$ (6)
where the Bloch Hamiltonian
$H_{k}=\begin{pmatrix}0&(t+\Delta)&0&t{e^{-4ik}}\\ (t+\Delta)&0&t&0\\ 0&t&0&(t-\Delta)\\ t{e^{4ik}}&0&(t-\Delta)&0\end{pmatrix}.$ (7)
@@
${\epsilon(k)=\pm\sqrt{2t^{2}+\Delta^{2}\pm t\sqrt{2t^{2}+6\Delta^{2}+2(t^{2}-\Delta^{2})\cos 4k}}}$ (8)
It essentially gives four bands in the energy spectrum. Notice that the energy gap closes for $\Delta=0$ at $k=0$ as well as for $(\Delta/t)^{2}=2$ at $k=\pm\pi/4$. Of these, the 1st case is a result of using periodic boundary implied in our Fourier construction but the system gaps out there in a finite chain. However the 2nd case indicates a gapless point even for the finite chain and correspond to the gap closing topological phase transition point. The bands cross linearly there (see Fig.3) making the low energy modes behave as Dirac fermions[42].
The numerical energy spectra for SSH(like) chain under OBC for $\theta=\pi/2$ is shown in Fig.2. The figure indicates that energy spectra is symmetric (up-down) i.e. for every eigenstate with energy $E$, its chiral symmetric partner having energy $-E$ necessarily exists. Here, we get a 4 sublattice configuration yielding 2 in-gap non-zero energy states and 2 ZES. Zero energy modes with $L=256$ will no longer subsist for $-0.3\lesssim\Delta/t\lesssim 0.3$ and for $\Delta/t\gtrsim 1.4$ or $\Delta/t\gtrsim-1.4$ [see e.g. Figs.2(b)-(c)].
For an odd number of lattice sites, one gets odd number of zero energy states. We examined that in a chain with odd number of sites, there are $L~{}modulo~{}4$ number of extra ZES that are localized exactly at $Mod(L,4)$ end sites. For example for $L=255$, with spectra as shown in Fig.2(d), there are 5 (for $|\Delta/t|\gtrsim 0.25$) or 3 (for $|\Delta/t|\lesssim 0.25$) zero energy modes ($i.e.$, 3 extra ZES) with the extra modes being localized at site no. 253, 254 and 255 (it could have been at site number 1, 2, 3 on the other side of the chain had the Hamiltonian matrix taken the three extra bonds on that side).
Refer to captionRefer to captionRefer to caption(a)(b)(c)
Refer to caption(d)
Figure 2: (a) Spectra of a SSH(like) chain with $L=256$ considering OBC as a function of $\Delta/t$ for $\theta=\frac{\pi}{2}$. (b) and (c) shows the corresponding low energy states for $L=256$ with smaller and larger $\Delta/t$ respectively while (d) represents low energy spectra for an odd number of lattice sites with $L=255$.
Class T C S d=1 d=2 d=3
A 0 0 0 - Z -
AI +1 0 0 - - -
AII -1 0 0 - $\textbf{Z}_{2}$ $\textbf{Z}_{2}$
D 0 +1 0 $\textbf{Z}_{2}$ Z -
C 0 -1 0 - Z -
AIII 0 0 1 Z - Z
BDI +1 +1 1 Z - -
CI +1 -1 1 - - Z
DIII -1 +1 1 $\textbf{Z}_{2}$ $\textbf{Z}_{2}$ Z
CII -1 -1 1 Z - $\textbf{Z}_{2}$
@@ II.1.1 Symmetry and Winding Number
Now we discuss a bit more on the symmetry of the SSH(like) chain for $\theta=\frac{\pi}{2}$. By employing the chiral symmetric operator $\hat{S}=\Gamma_{4}=I_{2}\otimes\sigma_{z}$ with $I_{2}$ as the $2\times 2$ identity matrix and $\sigma$ indicating the Pauli matrices, one can check that $\{\Gamma_{4},H_{k}\}=0$. Hence, this Hamiltonian has a chiral symmetry (of course, for a chain with even number of sites). Apart from this, this Hamiltonian also respects the time-reversal symmetry ($T$) as well as particle-hole symmetry ($C$) and falls in the BDI class universality in a 10-fold way classification (see Table 1). Here $T$ ($C$) commutes (anti-commutes) with the Hamiltonian $H_{k}$ which, in this spinless case, is given by $\hat{T}=I_{2}\otimes I_{2}\mathcal{K}$ ($\hat{C}=I_{2}\otimes\sigma_{z}\mathcal{K}$), $\mathcal{K}$ denoting the complex-conjugation). The relation $\hat{S}=\hat{T}\hat{C}$ is thus satisfied[41]. Due to the chiral symmetry of Hamiltonian, the Bloch Hamiltonian [Eq.(7)] is expected to be off-diagonal. We should add here that the presence of chiral symmetry does not necessarily ensure $T$ and $C$ symmetry as there is a symmetry class called AIII that respects chiral symmetry alone. In fact, if we consider the hopping amplitudes to be complex, instead of real, then time-reversal symmetry no longer persists and $H_{SSH}$ moves on to the AIII universality class[45]. The AIII class, in one dimension ($1D$), is always topologically nontrivial[43], unlike the usual SSH chain (with $\theta=\pi$), and its topology is described by a topological winding number.
There are two important corollaries of this chiral symmetry for the eigenstates and energies of the Hamiltonian. Firstly, the energy spectrum of the SSH(like) chain is always symmetric (up-down symmetry) in nature. For each eigenstate $\psi_{k}$ having energy $E$, there is always another eigenstate (its chiral partner, $\Gamma_{4}\psi_{k}=\psi^{\prime}_{k}$) with energy $-E$. Secondly, states with zero energy only occupy a single sub-lattice. When the total number of particles is odd, chiral symmetry infers that there will be an odd number of end modes having zero energy. Moreover, the energy of the in-gap end states can also be non-zero, and this symmetry implies that such states having non-zero energy should be even in number and symmetrically positioned about $E=0$.
@@
$V(k)=\begin{pmatrix}(t+\Delta)&t{e^{-4ik}}\\ t&(t-\Delta)\end{pmatrix}$ (10)
@@
Now, the calculation of the winding number can be given by [46, 48]
$\mathcal{W}=-\frac{i}{2\pi}\int_{BZ}\partial_{k}i\phi(k)dk~{}.$ (12)
with integration on the reduced BZ, $k\in$ [$-\pi/4$$\pi/4$][51]. Here $\phi_{k}$ is the phase of the complex number ${\bf Det[V(k)]}=(t^{2}-\Delta^{2})-t^{2}e^{-4ik}=R(k)e^{i\phi(k)}$. For a two-band model (see e.g. Refs.[47, 52]) one does not need to consider the determinant as $V(k)$ is just a number. However, for systems with 4 bands or more, this formula aims to give the topological index without further diagonalizing the Hamiltonian Eq.(7)[53]. One can substitute $\ln{\bf Det}[V(k)]$ in place of $i\phi_{k}$ in Eq.(12)[46]. Now, the winding number Eq.(12), considering the determinant, takes the form of [43, 49, 55, 50, 54]
@@
and the estimate comes out to be
${\mathcal{W}=\Bigg{\{}\begin{matrix}1,&0<\Delta^{2}/t^{2}<2;\\ 0,&\Delta^{2}/t^{2}>2;\\ \text{undefined},&\Delta/t=0\end{matrix}}$ (14)
The complex variable ${\bf Det[V(k)]}$ is called the winding vector (it has a magnitude and phase angle in the complex plane as already mentioned) as its number of winding about the origin as $k$ varies in [$-\pi/4$, $\pi/4$] [51, 56] determines the winding number.
Eq.(14) implies that, one can get $\mathcal{W}=1$ for small $\Delta/t\neq 0$ indicating the non-trivial topological phases (NTP) there. But $\Delta/t=0$ indicates an undefined ${\mathcal{W}}$ and hence no gap closing Lifshitz quantum phase transition (QPT) point. On the other hand, ${\mathcal{W}}$ remains unity till $0<|\Delta/t|<\sqrt{2}$ and vanishes for $|\Delta/t|>\sqrt{2}$. Thus QPT occurs at $\Delta=\pm\sqrt{2}t$ in this case. We should add here that interchanging the Majorana operators $\gamma_{i,A}$, $\gamma_{i,B}$ in Hamiltonian Eq.(3) has no significant effect (other than an overall sign change) since our considered model is not including the next-nearest-neighbor hopping amplitudes and chemical potentials. Following the above procedure, we attain
$V^{\prime}(k)=\begin{pmatrix}(t+\Delta)&t\\ t{e^{4ik}}&(t-\Delta)\end{pmatrix}$ (15)
causing merely a sign change in the winding number: $\mathcal{W}\rightarrow-\mathcal{W}$.
It is worth mentioning that the sign changes in winding number can occur by relabelling sublattices $\{A,B,C,D\}\Leftrightarrow\{D,C,B,A\}$ (or, interchanging the Majorana operators). This can also be visualized as a ‘choice’ of chiral symmetry operators. In the former case it is $\hat{S}=\Gamma_{4}$ acting in the sublattice basis while in the latter it is $\hat{S}=-\Gamma_{4}$. Moreover, the sign of the winding number is related to the winding direction of winding vector Det[V(k)] which in turn leads to the type of (unpaired) Majorana fermion residing at the boundary. This change of sign makes no difference when one considers a single chain, however for multiple chains coupled to each other, such liberty no more persists as the relative sign of the individual chains becomes a relevant quantity of the system topology and Hamiltonian no more remains block off-diagonal. Detailed illustrations of the same can be found in Ref.[46].
@@
In spirit of Eq.(12), one obtain
$\mathcal{W}=\frac{1}{2\pi}[\phi(\pi/4)-\phi(-\pi/4)]=\frac{1}{2\pi}\Delta\phi$ (16)